# Joint Resource-Power Allocation and UE Rank Selection in Multi-User MIMO Systems with Linear Transceivers

### v1

@@ III The Joint Resource-power Allocation and UE Rank selection Problem
For casting the joint resource-power allocation and UE rank selection as an optimization problem, we first need to obtain a suitable differentiable function that well-approximates $I_{Q_{max}}(.)$ in (11), where $Q_{max}$ is the highest-order constellation used. Fig. 1 plots the SE as a function of SNR $\rho$ for four functions; $\mathop{\mathrm{log}_{2}}\left(1+\rho\right)$ is the Shannon capacity, $I_{1024}(\rho)$ refers to the empirically-obtained mutual information between the input and output of an AWGN channel with $1024$-QAM, “Empirical SE" refers to the empirically obtained set of values of SNRs required to achieve a codeword error rate (CER) of $10^{-3}$ in a single-input single-ouput (SISO) AWGN channel for each MCS level of the combined tables of [8, Tables 5.1.3.1-2, 5.1.3.1-4] using 5G NR LDPC codes of length $2880$. As evident from the figure, the Shannon capacity is an overly optimistic approximation for practical communication systems that use finite-sized constellations (e.g., $2^{m}$-QAM) while $I_{1024}(\rho)$ is still too optimistic for systems that use practical error-correcting codes. A good approximation needs to be conservative for lower MCS indices due to the fact that lower MCS levels serve UEs that have relatively low SINR, and $\mathop{\mathrm{log}_{2}}\left(1+0.5\rho\right)$ is one such logarithmic approximation that is closer to practical rates than the theoretical mutual information. We would like to emphasize that $\mathop{\mathrm{log}_{2}}\left(1+0.5\rho\right)$ is just an exemplary function that will be used in the rest of this paper, and better differentiable functions can be obtained by suitable optimization.
As mentioned in the previous section, the BS needs to have reasonably good channel estimates to perform resource-power allocation. These channel estimates are not available for each PRB, but for groups of PRBs which we call PRBG in this paper. Let $n_{PRB}$ denote the total number of PRBs and $n_{RBG}$ the total number of PRBGs in the system so that we have a channel estimate for every $n_{PRB}/n_{RBG}$ PRBs. Let $\delta_{i,g}\in\{0,1\}$ indicate whether PRBG $g$ is allocated to UE $i$ ($\delta_{i,g}=1$) or not ($\delta_{i,g}=0$). Note that unlike in 5G NR, we do not restrict the frequency allocation to be contiguous in this paper. For every notation used in Section II, we use the subscript $g$ instead of $(f,t)$ to denote that the notation applies to PRBG $g$ instead of RE $(f,t)$. With this, the BS uses the channel estimates to obtain estimates of the SINR $\rho_{i,j,g}$ for the $j^{th}$ symbol of UE $i$ on PRBG $i$ using (5) for the uplink and (10) for the downlink, where, for the latter, the BS can receive the overall interference-plus-noise variance feedback from each UE. Henceforth, we use the notation ${\cal G}\triangleq\left\{1,\cdots,n_{RBG}\right\}$, ${\mathcal{N}}_{UE}\triangleq\left\{1,\cdots,N_{UE}\right\}$.
@@
Here, the weights satisfy $a_{i}>0,\forall i$. In this paper, we take $a_{i}=1,\forall i\in{\mathcal{N}}_{UE}$ and $\alpha=1$, which corresponds to the geometric mean (GM) rate optimization or proportional fairness [12].
@@ III-A Uplink Joint Resource-power Allocation and UE Rank Selection
The BS obtains the channel estimates $\hat{\mathbf{H}}_{i,g}\in\mathbb{C}^{n_{B}\times n_{l}^{(i)}}$ corresponding to the first $n_{l}^{(i)}$ columns of $\mathbf{R}_{g}^{-1/2}\mathbf{H}_{i,f,t}\mathbf{W}_{i,f,t}$ of Section II-A for PRBG $g$ (note that the interference-plus-noise covariance is now calculated at the PRBG level and not at the RE level). Let $\hat{\mathbf{H}}_{g}\triangleq\left[\hat{\mathbf{H}}_{1,g},\cdots,\hat{\mathbf{H}}_{N_{UE},g}\right]\in\mathbb{C}^{n_{B}\times n_{L}}$ where $n_{L}=\sum_{i=1}^{N_{UE}}n_{l}^{(i)}$. Then, with transmit powers $p_{i,j,g}>0$, $j=1,\cdots,n_{l}^{(i)}$, for UE $i$ on PRBG $g$, we get an SINR of $\rho_{i,j,g}=\lambda_{i,j,g}p_{i,j,g}$ for a hypothetically transmitted $j^{th}$ layer of UE $i$ on PRBG $g$, where $\lambda_{i,j,g}$ is the reciprocal of the $(\sum_{i^{\prime}=1}^{i-1}n_{l}^{(i^{\prime})}+j)^{th}$ diagonal element of $\left[\hat{\mathbf{H}}_{g}^{H}\hat{\mathbf{H}}_{g}\right]^{-1}$. Therefore, from (12) and the log approximation function $\mathop{\mathrm{log}_{2}}\left(1+0.5\rho\right)$, the normalized expected rate in bits per slot (normalized by $T$ times the number of subcarriers in a PRBG) for UE $i$ is given by
@@
Let $\rho_{max}$ correspond to the SNR required to achieve a SE of $r_{max}$, i.e., $\rho_{max}=2\left(2^{r_{max}}-1\right)$. With this, the goal of the resource allocation problem is to maximize the GM of the UE rates or equivalently, $\sum_{i=1}^{N_{UE}}\mathop{\mathrm{ln}}\left(R_{UE,i}\right)$, as follows:
$P_{UL}:\quad$ $\underset{p_{i,j,g},\delta_{i,g},n_{l}^{(i)}}{\max}$ $\sum_{i=1}^{N_{UE}}\mathop{\mathrm{ln}}\left(\sum_{g=1}^{n_{RBG}}\sum_{j=1}^{n_{l}^{(i)}}\mathop{\mathrm{log}_{2}}\left(1+0.5\delta_{i,g}\lambda_{i,j,g}p_{i,j,g}\right)\right)$ (15a)
subject to $\delta_{i,g}\in\{0,1\},\sum_{g\in{\cal G}}\delta_{i,g}\geq 1,n_{l}^{(i)}\leq n_{U}^{(i)},\forall i\in{\mathcal{N}}_{UE},\forall g\in{\cal G},$ (15b)
$0\leq p_{i,j,g}\leq{\rho_{max}}/{\lambda_{{i,j,g}}},\forall i\in{\mathcal{N}}_{UE},\forall g\in{\cal G},j=1,\cdots,n_{l}^{(i)},$ (15c)
$\sum_{g\in{\cal G}}\sum_{j=1}^{n_{U}^{(i)}}\mathop{\mathrm{log}_{2}}\left(1+0.5\delta_{i,g}\lambda_{i,j,g}p_{i,j,g}\right)\geq r_{min}n_{l}^{(i)}\sum_{g\in{\cal G}}\delta_{i,g},\forall i\in{\mathcal{N}}_{UE},$ (15d)
$\sum_{g\in{\cal G}}\sum_{j=1}^{n_{l}^{(i)}}\delta_{i,g}p_{i,j,g}\leq P_{U,max},\forall i\in{\mathcal{N}}_{UE}.$ (15e)
In the above optimization problem, (15b) imposes an allocation requirement of at least one PRBG for each UE but any arbitrary value $n_{RBG,min}\geq 1$ is possible. Also, (15c) can be replaced by a UE-specific maximum power that would cater to power allocation for UEs with different quality-of-services (QoSs), or to limit the ICI. In the latter case, $\rho_{max}$ is the equivalent of $(P_{0},\alpha)$ for uplink power control (see Section V-A). The constraint in (15d) is a reformulation of (12).
@@ III-B Downlink Joint Resource-power Allocation and UE Rank Selection
The setup is as described in Section II-B with the usage of notation and the PRBG indices in the subscripts of symbols as explained in Section III-A. Without loss of generality, we assume that $\sigma_{i}^{2}=1$, $\forall i=1,\cdots,N_{UE}$, so that from (10), the normalized expected rate in bits per slot (normalized by $T$ times the number of subcarriers in a PRBG) for UE $i$ is given by
$R_{UE,i}=\sum_{g=1}^{n_{RBG}}\sum_{j=1}^{n_{l}^{(i)}}\mathop{\mathrm{log}_{2}}\left(1+\delta_{i,g}\lambda_{i,j,g}p_{i,j,g}\right).$ (16)
The resource allocation problem can be expressed as follows:
$P_{DL}:\quad$ $\underset{p_{i,j,g},\delta_{i,g},n_{l}^{(i)}}{\max}$ $\sum_{i=1}^{N_{UE}}\mathop{\mathrm{ln}}\left(\sum_{g=1}^{n_{RBG}}\sum_{j=1}^{n_{l}^{(i)}}\mathop{\mathrm{log}_{2}}\left(1+0.5\delta_{i,g}\lambda_{i,j,g}p_{i,j,g}\right)\right)$ (17a)
subject to $\delta_{i,g}\in\{0,1\},\sum_{g\in{\cal G}}\delta_{i,g}\geq 1,n_{l}^{(i)}\leq n_{U}^{(i)},\forall i\in{\mathcal{N}}_{UE},\forall g\in{\cal G},$ (17b)
$0\leq p_{i,j,g}\leq{\rho_{max}}/{\lambda_{{i,j,g}}},\forall i,\forall g\in{\cal G},j=1,\cdots,n_{l}^{(i)},$ (17c)
$\sum_{g\in{\cal G}}\sum_{j=1}^{n_{U}^{(i)}}\mathop{\mathrm{log}_{2}}\left(1+0.5\delta_{i,g}\lambda_{i,j,g}p_{i,j,g}\right)\geq r_{min}n_{l}^{(i)}\sum_{g\in{\cal G}}\delta_{i,g},\forall i\in{\mathcal{N}}_{UE},$ (17d)
$\sum_{k=1}^{n_{B}}\sum_{i=1}^{N_{UE}}\sum_{j=1}^{n_{l}^{(i)}}\sum_{g\in{\cal G}}\delta_{i,g}a_{k,i,j,g}p_{i,j,f,t}\leq P_{B,max},$ (17e)
$\sum_{i=1}^{N_{UE}}\sum_{j=1}^{n_{l}^{(i)}}\sum_{g\in{\cal G}}\delta_{i,g}a_{k,i,j,g}p_{i,j,g}\leq P_{ant},\forall k=1,\cdots,n_{B}.$ (17f)
Except for the expressions for the post-equalization SINR, the only other significant difference between the problem formulations $P_{UL}$ and $P_{DL}$ is that for the downlink, the available BS transmit power needs to be shared by all the UE as shown in (17e), and that there is also an additional PAPC at the BS as shown in (17f).
@@ IV The Proposed Optimization Technique
SINR estimation (5) requires computationally intensive matrix inversions, potentially involving Moore-Penrose inversion, singular value decomposition (SVD) or eigenvalue decomposition (EVD) depending on the ZF-precoding type used [1]. Since $P_{UL}$ and $P_{DL}$ are both non-convex, there is no guarantee that any proposed approach might lead to an optimal solution. Hence, the goal is to obtain a possibly sub-optimal solution without incurring a high computational complexity. In this regard, we split both optimization problems into two convex optimization stages with an intermediate processing stage in between. In lieu of $P_{UL}$ and $P_{DL}$, we consider the following problems:
$\bar{P}_{UL}:\quad$ $\underset{p_{i,j,g}}{\min}$ $-\sum_{i=1}^{N_{UE}}\mathop{\mathrm{ln}}\left(\sum_{g=1}^{n_{RBG}}\sum_{j=1}^{n_{l}^{(i)}}\mathop{\mathrm{log}_{2}}\left(1+0.5\delta_{i,g}\lambda_{i,j,g}p_{i,j,g}\right)\right)$ (18a)
subject to $0\leq p_{i,j,g}\leq{\rho_{max}}/{\lambda_{{i,j,g}}},\forall i\in{\mathcal{N}}_{UE},\forall g\in{\cal G},j=1,\cdots,n_{l}^{(i)},$ (18b)
@@
$\bar{P}_{DL}:\quad$ $\underset{p_{i,j,g}}{\min}$ $-\sum_{i=1}^{N_{UE}}\mathop{\mathrm{ln}}\left(\sum_{g=1}^{n_{RBG}}\sum_{j=1}^{n_{l}^{(i)}}\mathop{\mathrm{log}_{2}}\left(1+0.5\delta_{i,g}\lambda_{i,j,g}p_{i,j,g}\right)\right)$ (19a)
subject to $0\leq p_{i,j,g}\leq{\rho_{max}}/{\lambda_{{i,j,g}}},\forall i\in{\mathcal{N}}_{UE},\forall g\in{\cal G},j=1,\cdots,n_{l}^{(i)},$ (19b)
@@
$\sum_{i=1}^{N_{UE}}\sum_{j=1}^{n_{l}^{(i)}}\sum_{g=1}^{n_{RBG}}\delta_{i,g}a_{k,i,j,g}p_{i,j,g}\leq P_{ant},\forall k=1,\cdots,n_{B},$ (19d)
for the downlink. Note that in both $\bar{P}_{UL}$ and $\bar{P}_{DL}$, the $\delta_{i,g}$ and $n_{l}^{(i)}$ are fixed and not optimized over. Further, the non-convex constraints (15d) and (17d) are not considered. With this, both $\bar{P}_{UL}$ and $\bar{P}_{DL}$ are now convex-optimization problems with convex objectives and affine constraints and can be solved using interior point methods [13] for which there exist efficient large-scale optimization software like Interior Point OPTimizer (IPOPT) [14]. Let $\textrm{OPT}\left(\bar{P}\left(\left\{\lambda_{i,j,g},\delta_{i,g},n_{l}^{(i)},\forall i,j,g\right\}\right)\right)$ denote the solution to the convex optimization problem $\bar{P}$, where $\bar{P}$ is either $\bar{P}_{UL}$ or $\bar{P}_{DL}$ as the case may be. Further, we denote by $f\left(\hat{\mathbf{H}}_{g},\left\{\delta_{i,g},n_{l}^{(i)},\forall i\in{\mathcal{N}}_{UE}\right\}\right)$ the set of values $\lambda_{i,j,g},\forall i,j,g$. Note that for the uplink (see Sec. III-A), $\lambda_{i,j,g}$ is the reciprocal of the $(\sum_{i^{\prime}=1}^{i-1}n_{l}^{(i^{\prime})}+j)^{th}$ diagonal element of $\left[\hat{\mathbf{H}}_{g}^{H}\hat{\mathbf{H}}_{g}\right]^{-1}$, while for the downlink, it is obtained from the product of the channel matrix and the ZF precoder (see Sec. II-B). In both cases, $\lambda_{i,j,g}$ depends on the set of UEs that share the PRBG and the number of layers for each such UE. Therefore, we add the following definition, $\forall i\in{\mathcal{N}}_{UE},\forall g\in{\cal G}$ :
@@
7:$\left\{p_{i,j,g}^{(1)}\forall i,j,g\right\}\leftarrow\textrm{OPT}\left(\bar{P}\left(\left\{\lambda_{i,j,g},\delta_{i,g},n_{l}^{(i)},\forall i,j,g\right\}\right)\right)$
@@
13:     $r_{i,sum}\leftarrow\sum_{g=1}^{n_{RGB}}\delta_{i,g}\sum_{j=1}^{n_{l}^{(i)}}r_{i,j,f}$, $r_{i,mean}\leftarrow{r_{i,sum}}/{n_{l}^{(i)}\sum_{g=1}^{n_{RGB}}\delta_{i,g}}$
@@
20:              ${\cal A}_{i}\leftarrow\arg\max_{{\cal A}\subset{\cal G},|{\cal A}|=n_{RBG,min}}\left\{\sum_{g\in{\cal A}}\sum_{j=1}^{n_{l}^{(i)}}r_{i,j,g}\right\}$
@@
28:$\left\{p_{i,j,g}^{*}\forall i,j,g\right\}\leftarrow\textrm{OPT}\left(\bar{P}\left(\left\{\lambda_{i,j,g},\delta_{i,g}^{*},n_{l}^{(i)},\forall i,j,g\right\}\right)\right)$
The details of the second stage are better explained using Algorithm 1. Steps 9–10 identify the weak PRBGs and layers for each UE. Next, the idea is to start conservatively at $n_{l}^{(i)}=1$ layer for each UE and check iteratively if additional layers can be accommodated. Steps 11–16 perform these operations, as the number of layers remains constant across all allocated PRBGs for each UE. Note that Step 13 is used to ensure that (15d) or (17d) is respected. Once the number of layers has been identified for each UE, the next check is to ensure that the minimum number of PRBGs $n_{RBG,min}$ is assigned to each UE. If not, Steps 20–21 assign the PRBGs with the $n_{RBG,min}$ highest values of estimated UE rates for the UE in context. Finally, once the PRBG allocation and UE ranks have been arrived at, the values of $\lambda_{i,j,g}$ are recalculated in Step 27 if there is a change in the PRBG allocation and the UE ranks with respect to Stage 1. Step 28 ensures that the PRBGs that are not allocated and the UE layers that are not selected will play no further part in the optimization. If each UE is servable (i.e., it is possible to guarantee a minimum rate for that UE for some feasible resource-power allocation), by choosing to maximize the GM rates and due to Steps 11-16, we would have already ensured that the finally allocated PRBGs and layers for each UE are such that each UE has a minimum rate guarantee in the slot. Next, with $\rho_{min}\triangleq 2\left(2^{r_{min}}-1\right)$, we set the lower limit on $p_{i,j,g}$ in (18b) and (19b) to $\rho_{min}/\lambda_{i,j,g}$ if $\delta_{i,g}^{*}=1$ and $j=1,\cdots,n_{l}^{(i)}$, and 00 otherwise. Then, the output of Step 27 provides the final allocated powers to each layer and each PRBG for each UE.

### v2

@@ III The Joint Resource-power Allocation and UE Rank selection Problem
For casting the joint resource-power allocation and UE rank selection as an optimization problem, we first need to obtain a suitable differentiable function that well-approximates $I_{Q_{max}}(.)$ in (11), where $Q_{max}$ is the highest-order constellation used. Fig. 1 plots the SE as a function of SNR, denoted by $\rho$, for four functions; $\mathop{\mathrm{log}_{2}}\left(1+\rho\right)$ is the Shannon capacity, $I_{1024}(\rho)$ refers to the empirically-obtained mutual information between the input and output of an AWGN channel with $1024$-QAM, “Empirical SE" refers to the empirically obtained set of values of SNRs required to achieve a codeword error rate (CER) of $10^{-3}$ in a single-input single-ouput (SISO) AWGN channel for each MCS level of the combined tables of [8, Tables 5.1.3.1-2, 5.1.3.1-4] using 5G NR LDPC codes of length $2880$. As evident from the figure, the Shannon capacity is an overly optimistic approximation for practical communication systems that use finite-sized constellations (e.g., $2^{m}$-QAM) while $I_{1024}(\rho)$ is still too optimistic for systems that use practical error-correcting codes. A good approximation needs to be conservative for lower MCS indices due to the fact that lower MCS levels serve UEs that have relatively low SINR, and $\mathop{\mathrm{log}_{2}}\left(1+0.5\rho\right)$ is one such logarithmic approximation that is closer to practical rates than the theoretical mutual information. We would like to emphasize that $\mathop{\mathrm{log}_{2}}\left(1+0.5\rho\right)$ is just an exemplary function that will be used in the rest of this paper, and better differentiable functions can be obtained by suitable optimization.
As mentioned in the previous section, the BS needs to have reasonably good channel estimates to perform resource-power allocation. These channel estimates are not available for each PRB, but for groups of PRBs which we call PRBG in this paper. Let $n_{PRB}$ denote the total number of PRBs and $n_{RBG}$ the total number of PRBGs in the system so that we have a channel estimate for every $n_{PRB}/n_{RBG}$ PRBs. Let $\delta_{i,g}\in\{0,1\}$ indicate whether PRBG $g$ is allocated to UE $i$ ($\delta_{i,g}=1$) or not ($\delta_{i,g}=0$). Note that unlike in 5G NR, we do not restrict the frequency allocation to be contiguous in this paper. For every notation used in Section II, we use the subscript $g$ instead of $(f,t)$ to denote that the notation applies to PRBG $g$ instead of RE $(f,t)$. With this, the BS uses the channel estimates to obtain estimates of the SINR $\rho_{i,j,g}$ for the $j^{th}$ symbol of UE $i$ on PRBG $i$ using (5) for the uplink and (10) for the downlink, where, for the latter, the BS uses the interference-plus-noise variance feedback from each UE. Henceforth, we use the notation ${\cal G}\triangleq\left\{1,\cdots,n_{RBG}\right\}$, ${\mathcal{N}}_{UE}\triangleq\left\{1,\cdots,N_{UE}\right\}$.
@@
Here, the weights satisfy $a_{i}>0,\forall i$. In this paper, we take $a_{i}=1,\forall i\in{\mathcal{N}}_{UE}$ and $\alpha=1$, which corresponds to the geometric mean (GM) rate optimization or proportional fairness [12]. However, our problem formulation is applicable to any value of $\alpha$.
@@ III-A Uplink Joint Resource-power Allocation and UE Rank Selection
The BS obtains the channel estimates $\hat{\mathbf{H}}_{i,g}\in\mathbb{C}^{n_{B}\times n_{l}^{(i)}}$ corresponding to the first $n_{l}^{(i)}$ columns of $\mathbf{R}_{g}^{-1/2}\mathbf{H}_{i,f,t}\mathbf{W}_{i,f,t}$ of Section II-A for PRBG $g$ (note that the interference-plus-noise covariance is now calculated at the PRBG level and not at the RE level). Let $\hat{\mathbf{H}}_{g}\triangleq\left[\hat{\mathbf{H}}_{1,g},\cdots,\hat{\mathbf{H}}_{N_{UE},g}\right]\in\mathbb{C}^{n_{B}\times n_{L}}$ where $n_{L}=\sum_{i=1}^{N_{UE}}n_{l}^{(i)}$. Then, with transmit powers $p_{i,j,g}>0$, $j=1,\cdots,n_{l}^{(i)}$, for UE $i$ on PRBG $g$, we get an SINR of $\rho_{i,j,g}=\lambda_{i,j,g}p_{i,j,g}$ for a hypothetically transmitted $j^{th}$ layer of UE $i$ on PRBG $g$, where $\lambda_{i,j,g}$ is the reciprocal of the $(\sum_{i^{\prime}=1}^{i-1}n_{l}^{(i^{\prime})}+j)^{th}$ diagonal element of $\left[\hat{\mathbf{H}}_{g}^{H}\hat{\mathbf{H}}_{g}\right]^{-1}$. Therefore, from the log approximation function $\mathop{\mathrm{log}_{2}}\left(1+0.5\rho\right)$, the normalized expected rate in bits per slot (normalized by $T$ times the number of subcarriers in a PRBG) for UE $i$ is given by
@@
Let $\rho_{max}$ correspond to the SNR required to achieve a SE of $r_{max}$, i.e., $\rho_{max}=2\left(2^{r_{max}}-1\right)$. With this, the goal of the resource allocation problem is to maximize the GM of the UE rates, or equivalently, $\sum_{i=1}^{N_{UE}}\mathop{\mathrm{ln}}\left(R_{UE,i}\right)$, as follows:
$P_{UL}:$
$\underset{p_{i,j,g},\delta_{i,g},n_{l}^{(i)}}{\max}$ $\sum_{i=1}^{N_{UE}}\mathop{\mathrm{ln}}\left(R_{UE,i}\right)$ (15a)
s.t. $\delta_{i,g}\in\{0,1\},\sum_{g\in{\cal G}}\delta_{i,g}\geq 1,\forall i,g,$ (15b)
$n_{l}^{(i)}\leq n_{U}^{(i)},\forall i,$ (15c)
$0\leq p_{i,j,g}\leq{\rho_{max}}/{\lambda_{{i,j,g}}},\forall i,\forall g,\forall j,$ (15d)
$R_{UE,i}\geq r_{min}n_{l}^{(i)}\sum_{g\in{\cal G}}\delta_{i,g},\forall i,$ (15e)
$\sum_{g\in{\cal G}}\sum_{j=1}^{n_{l}^{(i)}}\delta_{i,g}p_{i,j,g}\leq P_{U,max},\forall i.$ (15f)
In the above optimization problem, (15b) imposes an allocation requirement of at least one PRBG for each UE but any arbitrary value $n_{RBG,min}\geq 1$ is possible. Also, (15d) can be replaced by a UE-specific maximum power that would cater to power allocation for UEs with different quality-of-services (QoSs), or to limit the ICI. In the latter case, $\rho_{max}$ is the equivalent of $(P_{0},\alpha)$ for uplink power control (see Section V-A). The constraint in (15e) is a reformulation of (12).
@@ III-B Downlink Joint Resource-power Allocation and UE Rank Selection
The setup is as described in Section II-B with the usage of notation and the PRBG indices in the subscripts of symbols as explained in Section III-A. Without loss of generality, we assume that $\sigma_{i}^{2}=1$, $\forall i=1,\cdots,N_{UE}$, so that from (10), the normalized expected rate in bits per slot (normalized by $T$ times the number of subcarriers in a PRBG) for UE $i$ is the same as in (14). The resource allocation problem $P_{DL}$ can now be expressed as
$P_{DL}:$
$\underset{p_{i,j,g},\delta_{i,g},n_{l}^{(i)}}{\max}$ $\sum_{i=1}^{N_{UE}}\mathop{\mathrm{ln}}\left(R_{UE,i}\right)$ (16a)
s.t. $\delta_{i,g}\in\{0,1\},\sum_{g\in{\cal G}}\delta_{i,g}\geq 1,\forall i,\forall g,$ (16b)
$n_{l}^{(i)}\leq n_{U}^{(i)},\forall g,$ (16c)
$0\leq p_{i,j,g}\leq{\rho_{max}}/{\lambda_{{i,j,g}}},\forall i,\forall g,\forall j,$ (16d)
$R_{UE,i}\geq r_{min}n_{l}^{(i)}\sum_{g\in{\cal G}}\delta_{i,g},\forall i,$ (16e)
$\sum_{k=1}^{n_{B}}\sum_{i=1}^{N_{UE}}\sum_{j=1}^{n_{l}^{(i)}}\sum_{g\in{\cal G}}\delta_{i,g}a_{k,i,j,g}p_{i,j,f,t}\leq P_{B,max},$ (16f)
$\sum_{i=1}^{N_{UE}}\sum_{j=1}^{n_{l}^{(i)}}\sum_{g\in{\cal G}}\delta_{i,g}a_{k,i,j,g}p_{i,j,g}\leq P_{ant},\forall k.$ (16g)
Except for the expressions for the post-equalization SINR, the only other significant difference between the problem formulations $P_{UL}$ and $P_{DL}$ is that for the downlink, the available BS transmit power needs to be shared by all the UE as shown in (16f), and that there is also an additional PAPC at the BS as shown in (16g).
@@ IV The Proposed Optimization Technique
SINR estimation as given by (5) and (10) requires computationally intensive matrix inversions, potentially involving Moore-Penrose inversion, singular value decomposition (SVD) or eigenvalue decomposition (EVD) depending on the ZF-precoding type used [1]. Since $P_{UL}$ and $P_{DL}$ are both non-convex, there is no guarantee that any proposed approach might lead to an optimal solution. Hence, the goal is to obtain a possibly sub-optimal solution without incurring a high computational complexity. In this regard, we split both optimization problems into two convex optimization stages with an intermediate processing stage in between. In lieu of $P_{UL}$ and $P_{DL}$, we consider the following problems where $\delta_{i,g}$ and $n_{l}^{(i)}$ are fixed and not optimized over:
$\bar{P}_{UL}:$
$\underset{p_{i,j,g}}{\min}$ $-\sum_{i=1}^{N_{UE}}\mathop{\mathrm{ln}}\left(R_{UE,i}\right)$ (17a)
s.t. $0\leq p_{i,j,g}\leq{\rho_{max}}/{\lambda_{{i,j,g}}},\forall i,\forall g,\forall j,$ (17b)
@@
$\bar{P}_{DL}:$
$\underset{p_{i,j,g}}{\min}$ $-\sum_{i=1}^{N_{UE}}\mathop{\mathrm{ln}}\left(R_{UE,i}\right)$ (18a)
s.t. $0\leq p_{i,j,g}\leq{\rho_{max}}/{\lambda_{{i,j,g}}},\forall i,\forall g,\forall j,$ (18b)
@@
$\sum_{i=1}^{N_{UE}}\sum_{j=1}^{n_{l}^{(i)}}\sum_{g=1}^{n_{RBG}}\delta_{i,g}a_{k,i,j,g}p_{i,j,g}\leq P_{ant},\forall k,$ (18d)
for the downlink. Also to note is that the non-convex constraints (15e) and (16e) are not considered. With this, both $\bar{P}_{UL}$ and $\bar{P}_{DL}$ are now convex-optimization problems with convex objectives and affine constraints and can be solved using interior point methods [13] for which there exist efficient large-scale optimization software like Interior Point OPTimizer (IPOPT) [14]. Let $\textrm{OPT}\left(\bar{P}\left(\left\{\lambda_{i,j,g},\delta_{i,g},n_{l}^{(i)},\forall i,j,g\right\}\right)\right)$ denote the solution to the convex optimization problem $\bar{P}$, where $\bar{P}$ is either $\bar{P}_{UL}$ or $\bar{P}_{DL}$ as the case may be. Further, we denote by $f\left(\hat{\mathbf{H}}_{g},\left\{\delta_{i,g},n_{l}^{(i)},\forall i\in{\mathcal{N}}_{UE}\right\}\right)$ the set of values $\lambda_{i,j,g},\forall i,j,g$. Note that for the uplink (see Sec. III-A), $\lambda_{i,j,g}$ is the reciprocal of the $(\sum_{i^{\prime}=1}^{i-1}n_{l}^{(i^{\prime})}+j)^{th}$ diagonal element of $\left[\hat{\mathbf{H}}_{g}^{H}\hat{\mathbf{H}}_{g}\right]^{-1}$, while for the downlink, it is obtained from the product of the channel matrix and the ZF precoder (see Sec. II-B). In both cases, $\lambda_{i,j,g}$ depends on the set of UEs that share the PRBG and the number of layers for each such UE. Therefore, we add the following definition, $\forall i\in{\mathcal{N}}_{UE},\forall g\in{\cal G}$:
@@
7:$\left\{p_{i,j,g}^{(1)},\forall i,j,g\right\}\leftarrow\textrm{OPT}\left(\bar{P}\left(\left\{\lambda_{i,j,g},\delta_{i,g},n_{l}^{(i)},\forall i,j,g\right\}\right)\right)$
@@
13:     $r_{i,sum}\leftarrow\sum_{g=1}^{n_{RGB}}\delta_{i,g}\sum_{j=1}^{n_{l}^{(i)}}r_{i,j,f}$, $r_{i,mean}\leftarrow{r_{i,sum}}/\left({n_{l}^{(i)}\sum_{g=1}^{n_{RGB}}\delta_{i,g}}\right)$
@@
20:              ${\cal A}_{i}\leftarrow\underset{{\cal A}\subset{\cal G},|{\cal A}|=n_{RBG,min}}{\arg\max}\left\{\sum_{g\in{\cal A}}\sum_{j=1}^{n_{l}^{(i)}}r_{i,j,g}\right\}$
@@
28:$\left\{p_{i,j,g}^{*},\forall i,j,g\right\}\leftarrow\textrm{OPT}\left(\bar{P}\left(\left\{\lambda_{i,j,g},\delta_{i,g}^{*},n_{l}^{(i)},\forall i,j,g\right\}\right)\right)$
The details of the second stage are better explained using Algorithm 1. Steps 9–10 identify the weak PRBGs and layers for each UE. Next, the idea is to start conservatively at $n_{l}^{(i)}=1$ layer for each UE and check iteratively if additional layers can be accommodated. Steps 11–16 perform these operations, noting that the number of layers remains constant across all allocated PRBGs for each UE. Step 13 is used to ensure that (15e) or (16e) is respected. Once the number of layers has been identified for each UE, the next check is to ensure that the minimum number of PRBGs $n_{RBG,min}$ is assigned to each UE. If not, Steps 20–21 assign the PRBGs with the $n_{RBG,min}$ highest values of estimated UE rates for the UE in context. Finally, once the PRBG allocation and UE ranks have been arrived at, the values of $\lambda_{i,j,g}$ are recalculated in Step 27 if there is a change in the PRBG allocation and the UE ranks with respect to Stage 1, keeping also in mind (19) and (20). Step 28 ensures that the PRBGs that are not allocated and the UE layers that are not selected will play no further part in the optimization. If each UE is servable (i.e., it is possible to guarantee a minimum rate for that UE for some feasible resource-power allocation), by choosing to maximize the GM rates and due to Steps 11-16, we would have already ensured that the finally allocated PRBGs and layers for each UE are such that each UE has a minimum rate guarantee in the slot. Next, with $\rho_{min}\triangleq 2\left(2^{r_{min}}-1\right)$, we set the lower limit on $p_{i,j,g}$ in (17b) and (18b) to $\rho_{min}/\lambda_{i,j,g}$ if $\delta_{i,g}^{*}=1$ and $j=1,\cdots,n_{l}^{(i)}$, and 00 otherwise. Then, the output of Step 27 provides the final allocated powers to each layer and each PRBG for each UE.